# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ 6.3 Continuous Data: Lognormal-DDE for Response Times
Agree a lot 0.83 0.72 0.88 0.86 0.87
Agree a little 0.80 0.72 0.85 0.84 0.85
Disagree a little 0.76 0.74 0.85 0.84 0.84
Disagree a lot 0.69 0.67 0.76 0.75 0.78
@@
We fit the two-latent-layer DDE with $K^{(1)}=7,~{}K^{(2)}=1$, and estimate the latent skills based on the posterior probability using (14). We compare the estimated latent variables with the held-out information of each student’s categorical response to a survey question: “Mathematics is one of my favorite subjects”, and display the results in Table 5. The first column shows that the higher-order latent variable, $A_{1}^{(2)}$, is highly correlated with the magnitude that students like math. This suggests that $A_{1}^{(2)}$ can be interpreted as a general indicator for the students’ interest in math, while the fine-grained latent variables $\mathbf{A}^{(1)}$ represent more specific skill mastery for each domain. In addition, we observe that the students who enjoy math tend to have a higher probability of mastering specific skills as well. This is coherent with the fact that the estimated $\mathbf{B}^{(1)}$-coefficients are all positive.
@@ 7 Discussion
This paper makes contributions to core AI problems from statisticians’ perspectives by proposing a broad family of interpretable DGMs with solid identifiability guarantees, scalable computation pipelines, and promising application potential. It also opens up interesting directions for future research. Theoretically, it would be desirable to extend our identifiability and consistency results to the challenging setting with an unknown number of latent variables. We conjecture that one may be able to identify $\mathcal{K}$ under additional restrictions on the parameter space or considering specific parametric families. Second, it would be interesting to theoretically justify the good performance of the spectral-gap estimator for estimating the latent dimensions under general nonlinear link functions. Another interesting problem pertains to high-dimensional settings, where the number of observed responses, $J$, may grow with the sample size $N$. Our current notion of identifiability focuses on identifying the population model parameters under the traditional asymptotics with a fixed $J$, where the latent variables are marginalized out in the likelihood. As modern datasets often comes with a large number of observed features, it would be interesting to explore whether our identifiability and estimability results can be generalized to such settings.
In terms of the methodology and applications, an immediate future direction is working to effectively estimate and deploy deeper DDEs with $D\geq 3$ latent layers. The current theory and computational pipeline conceptually readily generalizes to deeper architectures than considered in this initial paper. It remains to test and refine these deeper models’ performance in larger-scale real-world data. Also, it would be interesting to extend DDEs to datasets with additional covariates. For example, the MNIST dataset comes with the actual digit labels as well as the spatial structure of the pixels in the image. In addition, many real datasets contain multiple modalities of data. For example, one may consider a dataset consisting of images uploaded in social media alongside text data such as tags. It is worthwhile to generalize DDEs to model such modern multimodal data. Finally, it would be interesting to extend DDEs for identifiable causal representation learning (Schölkopf et al.,, 2021) to uncover causal structures among the higher-order latent variables.
@@
Supplementary Material
This Supplementary Material is organized as follows. Section S.1 proves all main theorems and provides additional identifiability results under one-layer saturated models and variants. Section S.2 provides details regarding the spectral initialization algorithm. Section S.3 gives details regarding the EM algorithms (Algorithms 1 and 3) such as M-step update formulas, implementation details, and selection of the number of latent variables. Section S.4 provides various additional simulation results under (a) generically identifiable true parameters, (b) varying numbers of Monte Carlo samples in the SAEM algorithm, (c) unknown latent dimensions. Finally, Section S.5 gives additional data analysis details such as preprocessing, latent dimension selection, and additional visualizations.
@@
Recall from Section 3.1 that our identifiability results for general DDEs with multiple latent layers build upon the identifiability of a model with only one latent layer (the shallowest latent layer), where the deeper latent layers have been marginalized out. Here, we formally state identifiability conditions for such one-latent-layer saturated models in Section S.1.1, before proving the identifiability results for general DDEs in Section S.1.2. We prove the claims stated for the one-latent-layer models in Section S.1.3, and Theorem 3 in Section S.1.4. Additional identifiability results under related models with interaction effects (instead of the main-effect DDEs introduced in the main paper) are presented in Section S.1.5.
@@
For the one-latent-layer saturated model, define an equivalence relationship “$\sim_{K^{(1)}}$” by setting $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{K^{(1)}}(\widetilde{\boldsymbol{\Theta}}^{(1)},\widetilde{\mathbf{G}}^{(1)})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exist a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$ such that the following conditions hold:
@@
Consequently, the parameters $\mathbf{B}^{(d+1)},\mathbf{G}^{(d+1)},\boldsymbol{\pi}^{(d+1)}$ between the $d$th and $(d+1)$th layers are identified up to $\sim_{K^{(d+1)}}$. In particular, when $d=D-1$, it remains to determine ${\boldsymbol{p}}$ from the unstructured proportion parameter vector $\boldsymbol{\pi}^{(D)}$. Since we already have identified the $D$th layer labels up to $\sim_{K^{(D)}}$, this simply follows by marginalizing out the irrelevant coordinates:
@@
Now, assuming $\mathbf{B}^{(1)}\not\in N^{(1)}$, we can use a similar argument for deeper layers inductively. For $12\leq d\leq D-$ 1, let $N^{(d)}$ be the non-identifiable measure-zero subset of the $d$th layer coefficient space $\Omega(\mathbf{B}^{(d)};\mathbf{G}^{(d)})$. Note that we can still apply Proposition 2 since the conditional distribution of $\mathbf{A}^{(d)}\mid\mathbf{A}^{(d+1)}$ is modeled as a Bernoulli distribution with a logistic link $g_{\text{logistic}}$, which is indeed analytic and satisfy Assumption 2. Consequently, as long as $\mathbf{B}^{(d)}\not\in N^{(d)}$ for all $11\leq d\leq D-$ 1, the $D$-layer DDE is identifiable up to permutations of the latent variables within each layer. The proof is complete since $\cup_{d=1}^{D}(N^{(d)})^{c}$ is a union of a finite number of measure-zero sets, and hence again measure-zero. ∎
@@
We introduce additional notations that will be used to prove identifiability results for the one-latent-layer saturated model. Most of these notations are consistent with Lee and Gu, (2024). First, we omit the superscript “${(1)}$” that indicates the first latent layer when dealing with the one-latent-layer saturated model. Let $j,k,\boldsymbol{\alpha}$ denote typical indices for $j\in[J],k\in[K],\boldsymbol{\alpha}\in\{0,1\}^{K}$. Write $\mathbf{G}=\mathbf{G}^{(1)}=\{\mathbf{g}_{1}^{\top},\ldots,\mathbf{g}_{J}^{\top}\}^{\top}$ and let $H_{j}:=\{k\in[K]:G_{j,k}=1\}$ be the index of the parent latent variables for $Y_{j}$. Recall that for each $j$, the sample space $\mathcal{Y}_{j}$ is a separable metric space. Let $m_{j}$ be a base measure on $\mathcal{Y}_{j}$, this will be the counting measure for discrete sample spaces and the Lebesgue measure for continuous cases. Given $j$ and $\boldsymbol{\alpha}$, define the measure $\mathbb{P}_{j,\boldsymbol{\alpha}}$ on $\mathcal{Y}_{j}$ by setting
[…]
@@
$[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]:=\sum_{l=1}^{r}\mathbf{m}_{1,l}\circ\mathbf{m}_{2,l}\circ\mathbf{m}_{3,l}.$
@@
Consider the one-latent-layer saturated model, where the true parameters satisfy conditions A and B. Let $\mathbf{P}_{0}$ be the 3-way marginal probability tensor under these parameters. Then, the tensor decomposition $\mathbf{P}_{0}=[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]$ is unique up to a common column permutation. Here, $\mathbf{M}_{a}$s are $\upsilon_{a}\times 2^{K}$ matrices, whose last rows are the all-one vector $\mathbf{1}_{2^{K}}^{\top}$ for $a=2,3$. Additionally, assuming that the graphical matrix $\mathbf{G}$ is known, the model parameters $(\boldsymbol{\Theta},\mathbf{G})$ are identifiable up to sign-flipping.
@@
The following Lemma will be crucially utilized to identify and recover $\widetilde{\mathbf{G}}$. We present its proof after proving the main Proposition.
@@
Our proof builds upon Lemma S.1, which proved a more general notion of nonparametric identifiability of the one-layer model, but under a given graphical matrix $\mathbf{G}$. The cited Theorem 1 proves that the continuous parameters are identifiable up to sign flipping, given $\mathbf{G}$. In our setting, Assumption 1(c) resolves the sign flipping issue and the continuous parameters can be uniquely determined. Consequently, it suffices to show that $\mathbf{G}$ is identifiable up to the equivalence relation $\underset{K^{(1)}}{\sim}$. We separate this proof into two steps.
Step 1: Tensor decomposition and setup. Consider a one-latent-layer saturated model with true parameters $(\boldsymbol{\Theta},\mathbf{G})$ that satisfies conditions A, B. Suppose there exists an alternative set of parameters $(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ that define a same marginal distribution of $\mathbf{Y}$. Define the notation ~$\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}$ similar as ${\mathbb{P}}_{j,{\boldsymbol{\alpha}}}$ in (S.3), and also define $\widetilde{\mathbf{N}}_{1},\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ as the conditional probability matrices that specify the value of
@@
By applying Lemma S.1, the tensor decomposition in (S.6) is unique, so $\widetilde{\mathbf{N}}_{1}\text{{diag}}({\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}}),\mathbf{N}_{2},\mathbf{N}_{3}$ are identical up to a common column permutation, say $\mathfrak{S}\in S_{\{0,1\}^{K}}$. In particular, as the last row of $\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}})$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}})$ is exactly $\widetilde{\boldsymbol{\pi}}^{\top}$ and $\boldsymbol{\pi}^{\top}$, $\big{(}\widetilde{\boldsymbol{\pi}},\widetilde{\mathbf{N}}_{1}\big{)}$ and $\big{(}\boldsymbol{\pi},\mathbf{N}_{1}\big{)}$ are also identical up to the same permutation $\mathfrak{S}$.
@@
[…]
@@
Step 2: Proving the equivalence by constructing a column permutation. Now, we show that there exists a permutation $\sigma\in S_{[K]}$ such that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ for all $j\in[J],k\in[K]$. We first construct such a permutation $\sigma$ by observing $\mathcal{S}_{1},\ldots,\mathcal{S}_{K}$. For $k\in[K]$, (S.7) implies $|\mathcal{S}_{k}|=2$, and part (a) of Lemma S.2 constraints $\widetilde{\mathbf{g}}_{k}$ to be a standard basis vector. Hence, we can define $\sigma(k)$ such that $\widetilde{\mathbf{g}}_{k}=\mathbf{e}_{\sigma(k)}$. To see that $\sigma$ is indeed a permutation, we have to show that $\sigma(k)\neq\sigma(l)$ for $k\neq l$. But this is immediate by noting that the cardinality of the pmf vector of $\{\mathbb{P}(Y_{k}\in S_{1,k},Y_{l}\in S_{1,l}\mid\mathbf{A}=\boldsymbol{\alpha}):\boldsymbol{\alpha}\}$ is 2 if and only if $k=l$. For future purposes, let us partition the $2^{K}$ row indices $\widetilde{\boldsymbol{\alpha}}$ in (S.8) into two groups $T_{k}$ and $T_{k}^{c}$ based on the value of (~$\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}$. Then, the columns of $\widetilde{\mathbf{N}}$ that correspond to the $\widetilde{\boldsymbol{\alpha}}\in T_{k}$ must be identical to the columns of ${\mathbf{N}}$ indexed by $\boldsymbol{\alpha}$s such that $\alpha_{\sigma(k)}=0$ (or 1).
Now, for each $j>K$ and $k\in[K]$, we show that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and complete the proof. By part (b) of Lemma S.2, we have $\widetilde{g}_{j,k}$ = 0 if and only if
{(~$\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\in T_{k}\big{\}}=\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\not\in T_{k}\big{\}}.$ (S.9)
@@
Another application of part (b) of Lemma S.2 shows that this is equivalent to $g_{j,\sigma(k)}=0$. Hence, $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and the proof is complete. ∎
@@
    The proof is immediate from the faithfulness of the graphical matrix $\mathbf{G}$. For example, the third bullet point follows from noting that $|H_{j}|\geq 2$ implies |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\geq 3$, and that |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\leq 2$ implies $|H_{j}|\leq 1$.
@@
Our next lemma provides additional structure regarding the $2^{K}$ column indices in the tensor decomposition in Lemma S.3. This is a nontrivial problem as we wish to identify the parameters up to a label switching of $K$ binary latent variables. Here, to formally state the label switching, recall the notation of $\sim_{K}$ Definition S.2 and write $\mathbf{B}\sim_{K}\widetilde{\mathbf{B}},\mathbf{G}\sim_{K}\widetilde{\mathbf{G}}$, $\boldsymbol{\alpha}\sim_{K}\widetilde{\boldsymbol{\alpha}}$ when there exists a permutation $\sigma\in S_{[K]}$ such that $\beta_{j,l}=\widetilde{\beta}_{j,\sigma(l)}$, $g_{j,l}=\widetilde{g}_{j,\sigma(l)}$, and $\boldsymbol{\alpha}=(\widetilde{\alpha}_{\sigma(1)},\ldots,\widetilde{\alpha}_{\sigma(K)}).$
@@
Suppose that there exist two sets of parameters $(\mathbf{B},\mathbf{G})$, $(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}})$ that satisfies Assumption 1 and defines an identical $\eta_{j,\boldsymbol{\alpha}}$ in the following sense: there exists a permutation $\mathfrak{S}\in S_{\{0,1\}^{K}}$ such that
@@
We work under the same notations introduced at the beginning of the section. We make one additional assumption regarding the finite subsets $\mathcal{D}_{j}:=(S_{1,j},\ldots,S_{\kappa_{j},j})$ of the sample space $\mathcal{Y}_{j}$ as follows. We assume that $\mathcal{D}_{j}\subset\mathcal{C}_{j}$, where $\mathcal{C}_{j}$ is a countable separating class whose values determine probability measure on $\mathcal{Y}_{j}$. The existence of such a separating class is a consequence of $\mathcal{Y}_{j}$ being a separating metric space; see Step 1 in the proof of Theorem 1 in Lee and Gu, (2024) for a proof.
Suppose that $\mathbf{G}$ satisfy condition C, $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})$, and that there exists an alternate parameter $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$ that defines the same marginal likelihood. Here, we are dropping all superscripts in $\mathbf{G}=\mathbf{G}^{(1)\star}$ for simplicity. We show that $(\boldsymbol{\Theta},\mathbf{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ for $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$, where $\mathcal{N}_{1},\mathcal{N}_{2}$ are measure-zero sets that will be defined later. Recall that for $a=1,2,3$, $\mathbf{N}_{a}$ are the $\upsilon_{a}\times 2^{K}$ conditional probability matrices that describe $\mathbf{Y}\mid(\mathbf{A}=\boldsymbol{\alpha})$ under the true parameters $\boldsymbol{\Theta},\mathbf{G}$. Similarly, define $\widetilde{\mathbf{N}}_{a}$ as the corresponding matrices under the alternative parameters $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$. Similar to the proof of Proposition 1, we start from the rank $2^{K}$ decomposition of the marginal probability tensor in (S.5). By applying Lemma S.3, the decomposition

### v2

@@ 6.3 Multimodal Educational Data: Bernoulli-Lognormal-DDE
Agree a lot 0.62 0.59 0.62 0.42 0.57
Agree a little 0.50 0.47 0.52 0.34 0.43
Disagree a little 0.37 0.33 0.38 0.31 0.37
Disagree a lot 0.29 0.29 0.32 0.22 0.29
@@
We fit the two-latent-layer DDE with $K^{(1)}=7,~K^{(2)}=1$, and estimate the latent skills based on the posterior probability using (10) (see Supplement S.5.2 for details). We compare the estimated latent variables with the held-out information of each student’s categorical response to a survey question: “Mathematics is one of my favorite subjects”, and display the results in Table 6. The first column shows that the higher-order latent variable, $A_{1}^{(2)}$, is highly correlated with the extent that students like math. This suggests that $A_{1}^{(2)}$ can be interpreted as a general indicator of the students’ interest in math, while the fine-grained latent variables $\mathbf{A}^{(1)}$ represent students’ specific skill mastery profiles. In addition, we observe that the students who enjoy math tend to have a higher probability of mastering specific skills as well. This is coherent with the fact that the estimated $\mathbf{B}^{(1)}$-coefficients are nonnegative for both modes. In Supplement S.5.5, we also illustrate a strong correlation of the estimated intercepts for both data modalities, compared to another held-out information (whether or not the items are multiple choice questions).
@@ 7 Discussion
This paper makes significant contributions to core AI problems from statisticians’ perspective by proposing a broad family of interpretable DGMs with solid identifiability guarantees, scalable computational pipelines, and promising application potential. It also opens up interesting directions for future research. First, the current formulation of DDEs mainly focuses on binary latent variables and a multi-layer graphical structure, but it provides foundations for understanding more complex discrete latent variable models. We believe both the identifiability theory and estimation methods are extendable to general categorical/polytomous latent variables. Additionally, our theoretical identifiability guarantees extend to “generalized DDEs” with cross-level edges, which we illustrate in Supplement S.1.5. It would be interesting to propose suitable estimation methods for these more flexible model settings.
Another interesting problem pertains to high-dimensional settings, where the number of observed responses, $J$, may grow with the sample size $N$. Our current notion of identifiability focuses on identifying the population model parameters under the traditional asymptotics with a fixed $J$, where the latent variables are marginalized out in the likelihood. As modern datasets often comes with a large number of observed features, it would be interesting to explore whether our identifiability and estimability results can be generalized to such settings.
In terms of the methodology and applications, it would be interesting to extend DDEs to datasets with additional covariates. For example, the MNIST dataset comes with the actual digit labels as well as the spatial structure of the pixels in the image. Finally, it would be interesting to extend DDEs for identifiable causal representation learning (Schölkopf et al., , 2021) to uncover causal structures among the higher-order latent variables.
@@
Acknowldgement.
The authors are partially supported by NSF Grant DMS-2210796.
Supplement to “Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers”
This Supplementary Material is organized as follows. Section S.1 proves all main theorems and provides additional identifiability results under one-layer saturated models and generalized DDEs that goes beyond multi-layer structures. Section S.2 provides details regarding the spectral initialization algorithm. Section S.3 gives details regarding the EM algorithm) such as the penalized EM algorithm, M-step update formulas, implementation details, and selection of the number of latent variables. Section S.4 provides various additional simulation results under (a) generically identifiable true parameters, (b) varying numbers of Monte Carlo samples in the SAEM algorithm, (c) unknown latent dimensions. Section S.5 gives additional data analysis details such as preprocessing, latent dimension selection, and additional visualizations. Finally, Section S.6 discusses additional related works.
@@
Recall from Section 3.1 that our identifiability results for general DDEs with multiple latent layers build upon the identifiability of a model with only one latent layer (the shallowest latent layer), where the deeper latent layers have been marginalized out. Here, we formally state identifiability conditions for such one-latent-layer saturated models in Section S.1.1, before proving the identifiability results for general DDEs in Section S.1.2. We prove the claims stated for the one-latent-layer models in Section S.1.3, and Theorem 3 in Section S.1.4. Section S.1.5 establishes identifiability for generalized DDEs that allow additional edges compared. Additional identifiability results for selecting the latent dimension and results for related models with interaction effects (instead of the main-effect DDEs introduced in the main paper) are presented in Section S.1.6.
@@
For the one-latent-layer saturated model, define an equivalence relationship $\sim_{K^{(1)}}$” by setting $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{K^{(1)}}(\widetilde{\boldsymbol{\Theta}}^{(1)},\widetilde{\mathbf{G}}^{(1)})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exist a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$ such that the following conditions hold:
@@
Consequently, the parameters $\mathbf{B}^{(d+1)},\mathbf{G}^{(d+1)},\boldsymbol{\pi}^{(d+1)}$ between the $d$th and $(d+1)$th layers are identified up to $\sim_{K^{(d+1)}}$. In particular, when d=D1d=D-1, it remains to determine $\boldsymbol{p}$} from the unstructured proportion parameter vector $\boldsymbol{\pi}^{(D)}$. Since we already have identified the $D$th layer labels up to $\sim_{K^{(D)}}$, this simply follows by marginalizing out the irrelevant coordinates:
@@
Now, assuming $\mathbf{B}^{(1)}\not\in N^{(1)}$, we can use a similar argument for deeper layers inductively. For $2\leq d\leq D-1$, let $N^{(d)}$ be the non-identifiable measure-zero subset of the $d$th layer coefficient space $\Omega(\mathbf{B}^{(d)};\mathbf{G}^{(d)})$. Note that we can still apply Proposition 2 since the conditional distribution of $\mathbf{A}^{(d)}\mid\mathbf{A}^{(d+1)}$ is modeled as a Bernoulli distribution with a logistic link $g_{\text{logistic}}$, which is indeed analytic and satisfy Assumption 2. Consequently, as long as $\mathbf{B}^{(d)}\not\in N^{(d)}$ for all $1\leq d\leq D-1$, the $D$-layer DDE is identifiable up to permutations of the latent variables within each layer. The proof is complete since $\cup_{d=1}^{D}(N^{(d)})^{c}$ is a union of a finite number of measure-zero sets, and hence again measure-zero. ∎
@@
We introduce additional notations that will be used to prove identifiability results for the one-latent-layer saturated model. Most of these notations are consistent with Lee and Gu, (2024). First, we omit the superscript “(1){(1)}” that indicates the first latent layer when dealing with the one-latent-layer saturated model. Let $j,k,\boldsymbol{\alpha}$ denote typical indices for $j\in[J],k\in[K],\boldsymbol{\alpha}\in\{0,1\}^{K}$. Write $\mathbf{G}=\mathbf{G}^{(1)}=\{\mathbf{g}_{1}^{\top},\ldots,\mathbf{g}_{J}^{\top}\}^{\top}$ and let $H_{j}:=\{k\in[K]:G_{j,k}=1\}$ be the index of the parent latent variables for $Y_{j}$. Recall that for each $j$, the sample space $\mathcal{Y}_{j}$ is a separable metric space. Let $m_{j}$ be a base measure on $\mathcal{Y}_{j}$, this will be the counting measure for discrete sample spaces and the Lebesgue measure for continuous cases. Given $j$ and $\boldsymbol{\alpha}$, define the measure $\mathbb{P}_{j,\boldsymbol{\alpha}}$ on $\mathcal{Y}_{j}$ by setting
$\mathbb{P}_{j,\boldsymbol{\alpha}}(S):=\mathbb{P}(Y_{j}\in S\mid\mathbf{A}=\boldsymbol{\alpha})=\int_{S}p_{j}(y;\beta_{j,0}+\sum_{k}\beta_{j,k}\alpha_{k},\gamma_{j})dm_{j}(y).$ (S.3)
@@
$\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]:=\sum_{l=1}^{r}\mathbf{m}_{1,l}\circ\mathbf{m}_{2,l}\circ\mathbf{m}_{3,l}.$
@@
Consider the one-latent-layer saturated model, where the true parameters satisfy conditions A and B. Let $\mathbf{P}_{0}$ be the 3-way marginal probability tensor under these parameters. Then, the tensor decomposition $\mathbf{P}_{0}=[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]$ is unique up to a common column permutation. Here, $\mathbf{M}_{a}$s are $\upsilon_{a}\times 2^{K}$ matrices, whose last rows are the all-one vector $\mathbf{1}_{2^{K}}^{\top}$ for $a=2,3$. Additionally, assuming that the graphical matrix $\mathbf{G}$ is known, the model parameters $\boldsymbol{\Theta}$ are identifiable up to sign-flipping.
While Lee and Gu, (2024) establishes identifiability of the conditional distributions $\{\mathbb{P}_{j,\boldsymbol{\alpha}}\}$ (see (S.3)), recovering $\mathbf{B},\boldsymbol{\gamma}$ from $\{\mathbb{P}_{j,\boldsymbol{\alpha}}\}$ is straightforward since each $\mathbb{P}_{j,\boldsymbol{\alpha}}$ is the distribution of an identifiable parametric family.
@@
The following Lemma will be crucially utilized to identify and recover $\widetilde{\mathbf{G}}$. Lemma S.2 generalizes the fact that when $Y_{j}$ is a pure child of $A_{k}^{(1)}$ (in other words $|H_{j}|=1$), we have
$|\mathcal{S}_{j}|=|\{\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}|=|\{\mathbb{P}(Y_{j}\mid A_{k}^{(1)}=\alpha_{k}^{(1)}):\alpha_{k}^{(1)}=0,1\}|=2.$
Thus, by partitioning the set of conditional distributions corresponding to $Y_{j}$ which are pure children, the binary indices $\{\boldsymbol{\alpha}^{(1)}:\alpha_{k}^{(1)}=1\}$ and $\widetilde{\mathbf{G}}$ can be recovered (up to the permutation $\sigma\in S_{[K^{(1)}]}$). We present its proof after proving the main Proposition.
@@
Our proof builds upon Lemma S.1, which proved a more general notion of nonparametric identifiability of the one-layer model, but under a given graphical matrix $\mathbf{G}$. Lemma S.1 proves that the continuous parameters are identifiable up to sign flipping, given $\mathbf{G}$. In our setting, Assumption 1(c) resolves the sign flipping issue and the continuous parameters can be uniquely determined. Consequently, it suffices to show that $\mathbf{G}$ is identifiable up to the equivalence relation $\underset{K^{(1)}}{\sim}$. We separate this proof into two steps.
[…]
@@
[…]
@@
$=$ $\Big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:{\widetilde{\boldsymbol{\alpha}}}\Big{\}}$, (S.8)
@@
Step 2: Proving the equivalence by constructing a column permutation. Now, we show that there exists a permutation $\sigma\in S_{[K]}$ such that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ for all $j\in[J],k\in[K]$. We first construct such a permutation $\sigma$ by observing $\mathcal{S}_{1},\ldots,\mathcal{S}_{K}$. For $k\in[K]$, (S.7) implies $|\mathcal{S}_{k}|=2$, and part (a) of Lemma S.2 constraints $\widetilde{\mathbf{g}}_{k}$ to be a standard basis vector. Hence, we can define $\sigma(k)$ such that $\widetilde{\mathbf{g}}_{k}=\mathbf{e}_{\sigma(k)}$. To see that $\sigma$ is indeed a permutation, we have to show that $\sigma(k)\neq\sigma(l)$ for $k\neq l$. But this is immediate by noting that the cardinality of the pmf vector of $\{\mathbb{P}(Y_{k}\in S_{1,k},Y_{l}\in S_{1,l}\mid\mathbf{A}=\boldsymbol{\alpha}):\boldsymbol{\alpha}\}$ is 2 if and only if $k=l$. For future purposes, let us partition the $2^{K}$ row indices $\widetilde{\boldsymbol{\alpha}}$ in (S.8) into two groups $T_{k}$ and $T_{k}^{c}$ based on the value of $\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}$. Then, the columns of $\widetilde{\mathbf{N}}$ that correspond to the $\widetilde{\boldsymbol{\alpha}}\in T_{k}$ must be identical to the columns of $\mathbf{N}$} indexed by $\boldsymbol{\alpha}$s such that $\alpha_{\sigma(k)}=0$ (or 1).
Now, for each $j>K$ and $k\in[K]$, we show that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$. By part (b) of Lemma S.2, we have $\widetilde{g}_{j,k}$ = 0 if and only if
$\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\in T_{k}\big{\}}=\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\not\in T_{k}\big{\}}.$ (S.9)
@@
Another application of part (b) of Lemma S.2 shows that this is equivalent to $g_{j,\sigma(k)}=0$. Hence, $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and we have shown $\mathbf{G}\underset{K^{(1)}}{\sim}\mathbf{G}$. This completes the proof. ∎
@@
    The proof is immediate from the faithfulness of the graphical matrix $\mathbf{G}$. For example, the third bullet point follows from noting that $|H_{j}|\geq 2$ implies $|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\geq 3$, and that $|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\leq 2$ implies $|H_{j}|\leq 1$.
@@
Our next lemma provides additional structure regarding the $2^{K}$ column indices in the tensor decomposition in Lemma S.3. This is a nontrivial problem as we wish to identify the parameters up to a label switching of $K$ binary latent variables. Note that we can no longer utilize the pure children to address this problem, which was the strategy under the setting of strict identifiability. Here, to formally state the label switching, recall the notation of $\sim_{K}$ Definition S.2 and write $\mathbf{B}\sim_{K}\widetilde{\mathbf{B}},\mathbf{G}\sim_{K}\widetilde{\mathbf{G}}$, $\boldsymbol{\alpha}\sim_{K}\widetilde{\boldsymbol{\alpha}}$ when there exists a permutation $\sigma\in S_{[K]}$ such that $\beta_{j,l}=\widetilde{\beta}_{j,\sigma(l)}$, $g_{j,l}=\widetilde{g}_{j,\sigma(l)}$, and $\boldsymbol{\alpha}=(\widetilde{\alpha}_{\sigma(1)},\ldots,\widetilde{\alpha}_{\sigma(K)}).$
@@
[…]
@@
We work under the same notations introduced at the beginning of the section. We make one additional assumption regarding the finite subsets $\mathcal{D}_{j}:=(S_{1,j},\ldots,S_{\kappa_{j},j})$ of the sample space $\mathcal{Y}_{j}$ as follows. We assume that $\mathcal{D}_{j}\subset\mathcal{C}_{j}$, where $\mathcal{C}_{j}$ is a countable separating class whose values determine probability measure on $\mathcal{Y}_{j}$. The existence of such a separating class is a consequence of $\mathcal{Y}_{j}$ being a separable metric space; see Step 1 in the proof of Theorem 1 in Lee and Gu, (2024) for a proof.
[…]